OpenStudy (melissa_something):

Solve using substitution: ♥︎ x+y=-3 y+5x+3

OpenStudy (melissa_something):

This is easy, but I get stuck on simple things... :/

OpenStudy (dumbcow):

what part do you get stuck on? can you solve one equation for y ?

OpenStudy (melissa_something):

I get stuck on the part where you have to put one equation in the other and solve. && What do you mean solve one equation for y? :( @dumbcow

OpenStudy (dumbcow):

ok, i mean for equation y +5x = 3 can you use algebra to isolate "y" y = something

OpenStudy (melissa_something):

OMG my bad it's supposed to say y=5x+3

OpenStudy (dumbcow):

oh ok , well then its already solved for you now to substitute, you will take everything on right side and replace it with the "y" in other equation 

OpenStudy (dumbcow):

Now you have x + 5x +3 = -3 combine terms 6x +3 = -3 6x = -6 x = -1

OpenStudy (melissa_something):

Thank you so much! I wasnt sure if i put parenthesis or not! (:

OpenStudy (dumbcow):

yw

OpenStudy (melissa_something):

Um, also it says "If the system is inconsistant or has dependent equations say so?

OpenStudy (dumbcow):